A 1.5 kg ball is thrown upward with a velocity of 6 m/s.

a. What is the kinetic energy of the ball as it is thrown upward?


1
Expert's answer
2021-02-03T02:47:06-0500

a)

"KE=\dfrac{1}{2}mv^2=\dfrac{1}{2}\cdot1.5\ kg\cdot(6\ \dfrac{m}{s})^2=27\ J."